Find the magnitude of the electrical force and the

gravitational force between the electron and the proton in the

atom. The average distance between the two particles is 0.53 A


1
Expert's answer
2021-04-23T11:01:42-0400

"F_e = \frac{q_eq_p}{4πε_0R^2} \\
= \frac{(1.6 \times 10^{-19})^2}{4π(8.85 \times 10^{-12})(0.53 \times 10^{-10})^2} \\
= \frac{2.56 \times 10^{-38}}{(1.112 \times 10^{-10})(0.2809 \times 10^{-20})} \\
= \frac{2.56 \times 10^{-38}}{3.1236 \times 10^{-31}} \\
= 8.2 \times 10^{-8} \;N"